OpenStudy (anonymous):

g(x) = arcsin(4x + 8) Find the domain of this function.

OpenStudy (jamesj):

First, what is the domain of the function f(x) = arcsin(x) ?

OpenStudy (anonymous):

I know that the domain of h(x) = arcsin is [-1,1]. So, if I were to change it to the domain of g(x), wouldn't I just take each of the end points of the domain and apply the following: g(x) = arcsin(4(x+2)) So, to -1 and 1 individually, move left by 2 units, and multiply by 1/4, because they are horizontally stretched by a factor of 1/4?

OpenStudy (jamesj):

So\[-1 \leq 4x + 8 \leq 1\]

OpenStudy (jamesj):

The solution of that relation is the domain of your function

OpenStudy (anonymous):

What? How did you get that?

OpenStudy (anonymous):

The domain of arcsin(x) = [-1,1]

OpenStudy (jamesj):

The domain of f(y) = arcsin(y) is y in [-1,1] i.e., \[-1 \leq y \leq 1\]

OpenStudy (anonymous):

I am confused as to what the y is doing ther.e

OpenStudy (jamesj):

Here y = y(x) = 4x + 8. Hence \[-1 \leq 4x + 8 \leq 1 \]

OpenStudy (anonymous):

So only the terms within the argument of the function affect its domain?

OpenStudy (anonymous):

y is the range x is the domain

OpenStudy (jamesj):

One way to think about this function g is that it is the composite of two functions: f(x) = arcsin(x) and h(x) = 4x + 8 You function g(x) = f(h(x))

OpenStudy (jamesj):

So we need the RANGE of h(x) to be in the DOMAIN of f(x). The domain of f(x) = arcsin(x) is as we've observed [-1,1]

OpenStudy (anonymous):

Ok, so then how do you settle with the domain.

OpenStudy (jamesj):

So we need the range of h(x) which is 4x + 8 for given x to satisfy that relation as well.

OpenStudy (jamesj):

Hence solve \[-1 \leq 4x + 8 \leq 1\] to find the x for which the range of h is in the domain f. I.e., the x which is the domain for g(x) = arcsin(4x+8)

OpenStudy (anonymous):

Exactly right.

OpenStudy (anonymous):

So you are saying that where in the inner function does it work if it is to be composed? In this case, the inner functions' values are restricted from -1 to 1 because of the outer function's domain restrictions?

OpenStudy (jamesj):

yes, more or less, yes.

OpenStudy (jamesj):

For example, if x = 1, then 4x + 8 =12 and we can't evaluate arcsin(4x+8)